# Rethinking On-policy Optimization for Query Augmentation

### v1

@@ Abstract
Recent advances in large language models (LLMs) have led to a surge of interest in query augmentation for information retrieval (IR). Two main approaches have emerged. The first prompts LLMs to generate answers or pseudo-documents that serve as new queries, relying purely on the model’s parametric knowledge or contextual information. The second applies reinforcement learning (RL) to fine-tune LLMs for query rewriting, directly optimizing retrieval metrics. While having respective advantages and limitations, the two approaches have not been compared under consistent experimental conditions. In this work, we present the first systematic comparison of prompting-based and RL-based query augmentation across diverse benchmarks, including evidence-seeking, ad hoc, and tool retrieval. Our key finding is that simple, training-free query augmentation often performs on par with, or even surpasses, more expensive RL-based counterparts, especially when using powerful LLMs. Motivated by this discovery, we introduce a novel hybrid method, On-policy Pseudo-document Query Expansion (OPQE), which, instead of rewriting a query, the LLM policy learns to generate a pseudo-document that maximizes retrieval performance, thus merging the flexibility and generative structure of prompting with the targeted optimization of RL. We show OPQE outperforms both standalone prompting and RL-based rewriting, demonstrating that a synergistic approach yields the best results. Our implementation is made available to facilitate reproducibility. [^1] https://github.com/BaleChen/RethinkQAug
Rethinking On-policy Optimization for Query Augmentation
Zhichao Xu1, Shengyao Zhuang2, Xueguang Ma3, Bingsen Chen4, Yijun Tian5, Fengran Mo6, Jie Cao7, Vivek Srikumar1 1University of Utah, 2University of Queensland, 3University of Waterloo, 4New York University, 5University of Notre Dame, 6Université de Montréal, 7University of Oklahoma zhichao.xu@utah.edu          svivek@cs.utah.edu
@@ 1 Introduction
In many real-world retrieval scenarios, direct modification of the retriever is not possible — for example, when relying on API-based services such as OpenAI’s text-embedding-3 or proprietary systems like PubMed [^2] https://pubmed.ncbi.nlm.nih.gov/. In such settings, improving retrieval effectiveness must come from the query side. One common approach is supervised fine-tuning (SFT), where a model is trained on human- or weakly-annotated pairs of original queries and rewritten (or clarified/expanded) queries Yu et al. (2020); Mo et al. (2023). However, creating or obtaining high-quality query rewrite training data is nontrivial: humans may disagree about what makes a query “better”, domain shifts can limit transfer, and annotated datasets are expensive to build Yu et al. (2020); Ma et al. (2023); Coelho et al. (2025). Moreover, for dense retrievers the notion of what constitutes a “good” query is less well understood given their black-box neural architectures Faggioli et al. (2023); Meng et al. (2023). Therefore, there has been growing research interest for unsupervised or weakly-supervised query augmentation method, compared to data-centric SFT Wu et al. (2022).
Large language models (LLMs) have been applied to query augmentation in two ways: (1) by rewriting queries via SFT using (query, rewritten query) pairs (or weak supervision) to improve query clarity or specificity; and (2) by rewriting or expanding queries in a zero- or few-shot manner, without additional training Li et al. (2025). Prompt-based query augmentation methods such as HyDE Gao et al. (2023) and Query2Doc Wang et al. (2023) fall in the latter category: expanding queries with pseudo documents generated from an LLM’s paramatric knowledge can significantly improve retrieval effectiveness. These zero- or few-shot approaches require no additional training and are simple to apply. However, prior works Abe et al. (2025); Lei et al. (2024) have questioned these methods’ robustness, as generating a fluent or seemingly relevant document does not always guarantee a high-quality query representation. LLMs’ verbose outputs can introduce spurious terms, which may harm performance, particularly for sparse retrievers like BM25 Robertson et al. (1995).
@@
Recent work such as DeepRetrieval Jiang et al. (2025) represents one instantiation of this idea. It combines reasoning-style generation, inspired by frameworks like ReAct Yao et al. (2023) and DeepSeek-R1 Guo et al. (2025), with retrieval-based rewards to train LLMs that produce both intermediate reasoning steps and final queries. While DeepRetrieval demonstrates encouraging results across tasks including literature search, IR benchmarks, and text-to-SQL Liu et al. (2025), the approach remains resource-intensive and dependent on relevance-labeled data. More broadly, the field still lacks a systematic comparison between prompting-based query augmentation methods and RL-based formulations, leaving open questions about their respective strengths, weaknesses, and applicability in real-world settings.
In this work, we address this gap through a rigorous evaluation and benchmarking. Our contributions are threefold:
  • We perform a controlled comparison of DeepRetrieval against a simple, zero-shot prompting baseline across diverse retrieval tasks, including classical benchmarks and the emerging agent tool retrieval scenario.
  • Our experiments reveal a surprising and critical finding: the simple, training-free prompting method often matches or surpasses the complex and expensive RL-trained method. This observation suggests that leveraging the rich, structured knowledge of a powerful LLM via pseudo-document generation is a strong and resource-efficient baseline.
  • Based on this insight, we propose a novel hybrid method, On-policy Pseudo-document Query Expansion (OPQE), that combines the strengths of both paradigms. Our rationale is that while prompting provides a better structure for the augmented query (a pseudo-document), RL provides a powerful optimization mechanism. Our proposed modification reframes the RL task: instead of training the policy to rewrite a query, we train it to generate a full pseudo-document and use retrieval metrics as the reward. This synergy allows the model to leverage the rich generative format of prompting while fine-tuning the output to maximize retrieval scores. Our experiments confirm that this hybrid approach consistently achieves the best overall performance, outperform both prompt-based expansion method and RL-based query rewriting method.
@@ 2 Related Works
Query Expansion and Reformulation. Classical query expansion techniques include pseudo-relevance feedback Rocchio (1971); Voorhees (1994), translation-based expansion Cao et al. (2009), and statistical term weighting Zhai and Lafferty (2001). These approaches improve recall by introducing additional surface-level terms, but their effectiveness is often limited by noise sensitivity, especially for sparse retrievers.
@@
More recently, LLMs have enabled zero- and few-shot query rewriting. Methods such as HyDE Gao et al. (2023) and Query2Doc Wang et al. (2023) prompt an LLM to generate pseudo-documents to be encoded by retrievers. Related approaches like ThinkQE Lei et al. (2025) further incorporate reasoning before query generation. These methods are attractive because they are training-free and are directly applicable to API-based retrievers. However, their outputs are not explicitly optimized for retrieval metrics, and spurious terms may harm sparse retrieval performance Abe et al. (2025); Lei et al. (2024). Refer to Li et al. (2025) for an in-depth survey.
@@
Learning to Rewrite Queries. Beyond static prompting, several works explore trainable models for query reformulation. A prominent approach is supervised fine-tuning (SFT), where pre-trained sequence-to-sequence models like T5 Raffel et al. (2020) are trained on datasets of original and rewritten queries Wu et al. (2022); Liu et al. (2021). This technique is widely used in conversational search to transform ambiguous, context-dependent user utterances into clear, self-contained queries suitable for standard retrieval systems. However, the creation of high-quality, human-labeled datasets can be a significant bottleneck, and human rewrites may not always be optimal for retrieval performance Wu et al. (2022); Ye et al. (2023). To address this weakness, recent efforts have turned to weakly supervised methods. One such strategy involves knowledge distillation, where a powerful large language model (LLM) acts as a “teacher” to generate a large synthetic dataset of rewritten queries. A smaller, more efficient "student" model is then fine-tuned on this data, learning the rewriting task without expensive human annotation Ye et al. (2023). This methodology has proven effective within the “Rewrite-Retrieve-Read” framework for retrieval-augmented generation, where the initial query is adapted to be more effective for the downstream retrieval and reading components Ma et al. (2023).
@@
RL for Query Optimization. A growing line of work applies RL to optimize queries directly against retrieval-based rewards. Unlike supervised reformulation, which learns from a fixed dataset, RL enables on-policy exploration, directly tying the training signal to retrieval effectiveness Wang et al. (2020); Wu et al. (2022). This is particularly valuable in dynamic contexts like conversational search, where an RL agent can be trained to reformulate queries to maximize the performance of a dense retriever Zhu et al. (2025). This method’s key advantage is its ability to optimize a rewriter for a black-box retrieval system, where the retriever’s parameters are inaccessible for fine-tuning. The rewriter is treated as a policy model, and it receives a reward based on the performance of the retriever, allowing it to adapt and generate more effective queries. While powerful, these methods can require substantial compute for exploration and often depend on carefully engineered reward functions to guide the learning process Sheng et al. (2025); Lambert et al. (2024).
@@ 3 Background
Prompt-based Query Augmentation. Query augmentation aims to rewrite or expand an original query to improve retrieval effectiveness. Sparse methods such as BM25 Robertson et al. (1995) benefit directly from term-level expansions, while dense retrievers, which embed queries and documents into a shared vector space, pose a greater challenge since it is unclear what constitutes an effective query embedding. Collecting human-labeled training data for this task is expensive and often infeasible.
To reduce the dependency on labeled data, recent methods exploit the zero- and few-shot generative capabilities of LLMs. A common approach is to prompt an LLM to generate a pseudo-document that might answer the query. This pseudo-document is then encoded by the retriever to perform retrieval. Such prompt-based pseudo-document expansion avoids the need for explicit query-rewrite supervision and has been shown to be effective across open-domain QA and web search benchmarks Gao et al. (2023); Wang et al. (2023). However, its success is tightly coupled to the scale and internal knowledge of the LLM, raising questions about how to enable smaller models to serve as effective query rewriters.
RL-based Query Augmentation. Another approach is to optimize query rewriting via RL. Here, the query rewriter is treated as a policy $\pi_{\theta}(q^{\prime}\mid q)$ that generates an augmented query $q^{\prime}$ given the original query $q$. The process can be formalized as a Markov Decision Process (MDP), where State $s_{t}$ denotes the current context, typically the original query $q$ and the intermediate reasoning steps. Action $a_{t}$ means generating the next token or sub-query in $q^{\prime}$. Policy $\pi_{\theta}(a_{t}\mid s_{t})$ denotes LLM parameterized by $\theta$. Reward $r(s_{t},a_{t})$ is derived from retrieval effectiveness of the completed query $q^{\prime}$, e.g., Recall@K, NDCG. The objective is to maximize the expected retrieval reward, with a KL-regularization term relative to a reference policy $\pi_{\text{ref}}$ to stabilize training and maintain generation fluency:
$J_{\text{KL}}(\theta)$ $=\mathbb{E}_{q^{\prime}\sim\pi_{\theta}(\cdot\mid q)}\Big[r_{\text{retrieval}}(q^{\prime})$
$\quad-\beta\,\text{KL}\big(\pi_{\theta}(\cdot\mid q)\,\|\,\pi_{\text{ref}}(\cdot\mid q)\big)\Big].$ (1)
@@
Reasoning-Style Query Generation. Recent works have demonstrated the benefit of generating structured outputs, separating intermediate reasoning from the final answer. Frameworks like ReAct Yao et al. (2023) and DeepSeek-R1 Guo et al. (2025) inspire approaches where the policy first generates reasoning steps $h_{1},\dots,h_{n}$ and then a final augmented query $q^{\prime}_{\text{final}}$. The retrieval reward is computed against $q^{\prime}_{\text{final}}$, while intermediate steps may be guided via auxiliary heuristics or shaping rewards. DeepRetrieval Jiang et al. (2025) exemplifies this paradigm by combining reasoning-style generation with PPO training on retrieval-based rewards. While effective, this approach highlights general challenges in RL-based query augmentation: reward sparsity, exploration vs. fluency trade-offs, and computational cost for large LLMs.
Terminology. We use the term RL broadly to refer to optimization of query-rewriting policies via retrieval-based rewards. Specifically, our experiments employ on-policy optimization methods such as PPO Schulman et al. (2017). While RL encompasses other paradigms (e.g., off-policy learning), we focus on on-policy methods here.
Algorithm 1 Query Augmentation with PPO
1:Initialize LLM policy $\pi_{\theta}$ and reference policy $\pi_{\text{ref}}$
2:for each query $q$ in dataset do
3:  Generate augmented query $q^{\prime}\sim\pi_{\theta}(\cdot\mid q)$
4:  Compute retrieval reward $r_{\text{retrieval}}(q^{\prime})$
5:  Compute objective $J_{\text{KL}}(\theta)$
6:  Update $\theta$ via PPO
7:end for
Algorithm Overview. The procedure for RL-based query augmentation can be summarized as Algorithm˜1. This formulation allows training a query rewriter even when the retriever is a black-box system, such as an API-based service, and emphasizes the connection between query generation and retrieval performance.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have led to a surge of interest in query augmentation for information retrieval (IR). Two main approaches have emerged. The first prompts LLMs to generate answers or pseudo-documents that serve as new queries, relying purely on the model’s parametric knowledge or contextual information. The second applies reinforcement learning (RL) to fine-tune LLMs for query rewriting, directly optimizing retrieval metrics. While having respective advantages and limitations, the two approaches have not been compared under consistent experimental conditions. In this work, we present the first systematic comparison of prompting-based and RL-based query augmentation across diverse benchmarks, including evidence-seeking, ad hoc, and tool retrieval. Our key finding is that under a compute-aware comparison setting, simple, training-free query augmentation often performs on par with, or even surpasses, more expensive RL-based counterparts, especially when using powerful LLMs. Motivated by this discovery, we introduce a novel hybrid method, On-policy Pseudo-document Query Expansion (OPQE), which, instead of rewriting a query, the LLM policy learns to generate a pseudo-document that maximizes retrieval performance, thus merging the flexibility and generative structure of prompting with the targeted optimization of RL. We show OPQE outperforms both standalone prompting and RL-based rewriting, demonstrating that a synergistic approach yields the best results. We open source our implementation to facilitate reproducibility. [^1] https://github.com/BaleChen/RethinkQAug
@@ 1 Introduction
In many real-world retrieval scenarios, direct modification of the retriever is often not possible — for example, when relying on API-based services such as OpenAI’s text-embedding-3 or proprietary systems like PubMed [^2] https://pubmed.ncbi.nlm.nih.gov/. In such settings, improving retrieval effectiveness must come from the query side. One common approach is supervised fine-tuning (SFT), where a model is trained on human- or weakly-annotated pairs of original queries and rewritten (or clarified/expanded) queries (Yu et al., 2020; Mo et al., 2023). However, creating or obtaining high-quality query rewrite training data is nontrivial: humans may disagree about what makes a query “better”, domain shifts can limit transfer, and annotated datasets are expensive to build (Yu et al., 2020; Ma et al., 2023; Coelho et al., 2025). Moreover, for dense retrievers the notion of what constitutes a “good” query is less well understood given their black-box neural architectures (Faggioli et al., 2023; Meng et al., 2023). Therefore, there has been growing research interest in unsupervised or weakly-supervised query augmentation methods, compared to data-centric SFT (Wu et al., 2022).
Large language models (LLMs) have been applied to query augmentation in two ways: (1) by rewriting queries via SFT using (query, rewritten query) pairs (or weak supervision) to improve query clarity or specificity; and (2) by rewriting or expanding queries in a zero- or few-shot manner, without additional training (Li et al., 2025). Prompt-based query augmentation methods such as HyDE (Gao et al., 2023) and Query2Doc (Wang et al., 2023) fall in the latter category: expanding queries with pseudo-documents generated from an LLM’s parametric knowledge can significantly improve retrieval effectiveness. These zero- or few-shot approaches require no additional training and are simple to apply. However, prior works (Abe et al., 2025; Lei et al., 2024) have questioned these methods’ robustness, as generating a fluent or seemingly relevant document does not always guarantee a high-quality query representation. LLMs’ verbose outputs can introduce spurious terms, which may harm performance, particularly for sparse retrievers like BM25 (Robertson et al., 1995).
@@
Recent work such as DeepRetrieval (Jiang et al., 2025) represents one instantiation of this idea. It combines reasoning-style generation, inspired by frameworks like ReAct (Yao et al., 2023) and DeepSeek-R1 (Guo et al., 2025), with retrieval-based rewards to train LLMs that produce both intermediate reasoning steps and final queries. While DeepRetrieval demonstrates encouraging results across tasks including literature search, IR benchmarks, and text-to-SQL (Liu et al., 2025), the approach remains resource-intensive and dependent on relevance-labeled data. More broadly, the field still lacks a systematic comparison between prompting-based query augmentation methods and RL-based formulations, leaving open questions about their respective strengths, weaknesses, and applicability in real-world settings.
In this work, we address this gap through a rigorous evaluation and benchmarking. Therefore, we ask: under controlled and compute-aware comparisons, when does training-free pseudo-document generation suffice, and when does on-policy optimization provide meaningful benefits?
To answer this question, we run a controlled comparison between prompting-based pseudo-document query expansion and on-policy RL for direct query rewriting (DeepRetrieval’s formulation). Accounting for their different compute profiles — inference-time cost for prompting (which can be traded for stronger LLMs) versus training-time cost for RL — we find that vanilla on-policy RL for direct query rewriting often yields limited gains, and often outperformed by simple prompting. Moreover, RL effects are strongly task- and retriever-dependent: improvements are more consistent for sparse term-matching retrievers such as BM25, while gains for dense retrievers can be smaller or even negative depending on the task and reward design. These observations motivate On-policy Pseudo-document Query Expansion (OPQE), which applies on-policy optimization to pseudo-document generation rather than query rewriting, and yields more consistent improvements and the best overall performance.
Our contributions are threefold:
  • We perform a controlled comparison of DeepRetrieval against a simple, zero-shot prompting baseline across diverse retrieval tasks, including classical ad hoc retrieval and the emerging agentic tool retrieval.
  • Our experiments reveal a key finding: the simple, training-free prompting method often matches or surpasses more complex RL-trained methods. This observation suggests that leveraging the rich, structured knowledge of a powerful LLM via pseudo-document generation is a strong and resource-efficient baseline.
  • We propose OPQE, a hybrid method that combines the generative structure of prompting with the targeted optimization of on-policy RL by training the policy to generate pseudo-documents under retrieval-based rewards, and it yields the strongest overall retrieval performance across our benchmarks.
@@ 2 Related Works
Classical query expansion techniques include pseudo-relevance feedback (Rocchio, 1971; Voorhees, 1994), translation-based expansion (Cao et al., 2009), and statistical term weighting (Zhai & Lafferty, 2001). These approaches improve recall by introducing additional surface-level terms, but their effectiveness is often limited by noise sensitivity, especially for sparse retrievers.
@@
More recently, LLMs have enabled zero- and few-shot query rewriting. Methods such as HyDE (Gao et al., 2023) and Query2Doc (Wang et al., 2023) prompt an LLM to generate pseudo-documents to be encoded by retrievers. Related approaches like ThinkQE (Lei et al., 2025) further incorporate reasoning before query generation. These methods are attractive because they are training-free and are directly applicable to API-based retrievers. However, their outputs are not explicitly optimized for retrieval metrics, and spurious terms may harm sparse retrieval performance (Abe et al., 2025; Lei et al., 2024). Refer to Li et al. (2025) for an in-depth survey.
Learning to Rewrite Queries.
@@
Beyond static prompting, several works explore training models for query reformulation. A prominent approach is supervised fine-tuning, where pre-trained sequence-to-sequence models like T5 (Raffel et al., 2020) are trained on datasets of original and rewritten queries (Wu et al., 2022; Liu et al., 2021). This technique is widely used in conversational search to transform ambiguous, context-dependent user utterances into clear, self-contained queries suitable for standard retrieval systems. However, the creation of high-quality, human-labeled datasets can be a significant bottleneck, and human rewrites may not always be optimal for retrieval performance (Wu et al., 2022; Ye et al., 2023). To address this weakness, recent efforts have turned to weakly supervised methods. One such strategy involves knowledge distillation, where a powerful large language model (LLM) acts as a “teacher” to generate a large synthetic dataset of rewritten queries. A smaller, more efficient “student” model is then fine-tuned on this data, learning the rewriting task without expensive human annotation (Ye et al., 2023). This methodology has proven effective within the “Rewrite-Retrieve-Read” framework for retrieval-augmented generation, where the initial query is adapted to be more effective for the downstream retrieval and reading components (Ma et al., 2023).
RL for Query Optimization.
@@
A growing line of work applies RL to optimize queries directly against retrieval-based rewards. Unlike supervised reformulation, which learns from a fixed dataset that requires extensive annotation, RL enables on-policy exploration, directly tying the training signal to retrieval effectiveness (Wang et al., 2020; Wu et al., 2022) and requires no golden (query, rewritten query) pairs. This is particularly valuable in dynamic contexts like conversational search, where an RL agent can be trained to reformulate queries to maximize the performance of a dense retriever (Zhu et al., 2025). This method’s key advantage is its ability to optimize a rewriter for a black-box retrieval system, where the retriever’s parameters are inaccessible for fine-tuning. The rewriter is treated as a policy model, and it receives a reward based on the performance of the retriever, allowing it to adapt and generate more effective queries. While powerful, these methods can require substantial compute for exploration and often depend on carefully engineered reward functions to guide the learning process (Sheng et al., 2025; Lambert et al., 2024).
@@ 3 Background
Prompt-based Query Augmentation.
Query augmentation aims to rewrite or expand an original query to improve retrieval effectiveness. Sparse methods such as BM25 (Robertson et al., 1995) benefit directly from term-level expansions, while dense retrievers, which embed queries and documents into a shared vector space, pose a greater challenge since it is unclear what constitutes an effective query embedding. Collecting human-labeled training data for this task is expensive and often infeasible.
To reduce the dependency on labeled data, recent methods exploit the zero- and few-shot generative capabilities of LLMs. A common approach is to prompt an LLM to generate a pseudo-document that might answer the query. This pseudo-document is then encoded by the retriever to perform retrieval. Such prompt-based pseudo-document expansion avoids the need for explicit query-rewrite supervision and has been shown to be effective across open-domain QA and web search benchmarks (Gao et al., 2023; Wang et al., 2023). However, its success is tightly coupled to the scale and internal knowledge of the LLM, raising questions about how to enable smaller models to serve as effective query rewriters.
RL-based Query Augmentation.
Another approach is to optimize query rewriting via RL. Here, the query rewriter is treated as a policy $\pi_{\theta}(q^{\prime}\mid q)$ that generates an augmented query $q^{\prime}$ given the original query $q$. The process can be formalized as a Markov Decision Process (MDP), where State $s_{t}$ denotes the current context, typically the original query $q$ and the intermediate reasoning steps. Action $a_{t}$ means generating the next token or sub-query in $q^{\prime}$. Policy $\pi_{\theta}(a_{t}\mid s_{t})$ denotes LLM parameterized by $\theta$. Reward $r(s_{t},a_{t})$ is derived from retrieval effectiveness of the completed query $q^{\prime}$, e.g., Recall@K, NDCG. The objective is to maximize the expected retrieval reward, with a KL-regularization term relative to a reference policy $\pi_{\text{ref}}$ to stabilize training and maintain generation fluency:
$\mathcal{J}_{\text{KL}}(\theta)=\mathbb{E}_{q^{\prime}\sim\pi_{\theta}(\cdot\mid q)}\Big[r_{\text{retrieval}}(q^{\prime})-\beta\,\text{KL}\big(\pi_{\theta}(\cdot\mid q)\,\|\,\pi_{\text{ref}}(\cdot\mid q)\big)\Big].$ (1)
@@
Reasoning-Style Query Generation.
Recent works have demonstrated the benefit of generating structured outputs, separating intermediate reasoning from the final answer. Frameworks like ReAct (Yao et al., 2023) and DeepSeek-R1 (Guo et al., 2025) inspire approaches where the policy first generates reasoning steps $h_{1},\dots,h_{n}$ and then a final augmented query $q^{\prime}_{\text{final}}$. The retrieval reward is computed against $q^{\prime}_{\text{final}}$, while intermediate steps may be guided via auxiliary heuristics or shaping rewards. DeepRetrieval (Jiang et al., 2025) exemplifies this paradigm by combining reasoning-style generation with PPO training on retrieval-based rewards. While effective, this approach highlights general challenges in RL-based query augmentation: reward sparsity, exploration vs. fluency trade-offs, and computational cost for large LLMs.
Termininology.
We use the term RL broadly to refer to optimization of query-rewriting policies via retrieval-based rewards. Specifically, our experiments employ on-policy optimization methods such as PPO (Schulman et al., 2017). While RL encompasses other paradigms (e.g., off-policy learning), we focus on on-policy methods here.
Algorithm 1 Direct query rewriting with PPO
1:Initialize LLM policy $\pi_{\theta}$, reference policy $\pi_{\text{ref}}$, and value function $V_{\phi}$ 2:for each batch $\mathcal{B}$ from dataset do 3:  Initialize rollout buffer $\mathcal{D}\leftarrow\emptyset$ 4:  for each query $q\in\mathcal{B}$ do 5:    Sample a rollout (reasoning + augmented query): $\tau\sim\pi_{\theta}(\cdot\mid q)$ 6:    Extract augmented query $q^{\prime}$ from $\tau$ 7:    Retrieve top-$k$ passages with retriever 8:    Compute retrieval reward $r_{\text{retrieval}}(q^{\prime})$ with potential format constraints 9:    Assign the scalar reward to the rollout $\tau$ 10:    Estimate sequence-level advantages $\hat{A}(\tau)$ using $V_{\phi}$ (e.g., GAE) 11:    Add $(\tau,r_{\text{retrieval}}(q^{\prime}),\hat{A}(\tau))$ to $\mathcal{D}$ 12:  end for 13:  for each PPO epoch $e=1,\dots,E$ do 14:    for each minibatch $\mathcal{M}\subset\mathcal{D}$ do 15:     Update $\theta$ using PPO clipped objective + KL penalty to $\pi_{\text{ref}}$ 16:     Update $\phi$ by regression on returns (value loss) 17:    end for 18:  end for 19:end for
Algorithm Overview.
The procedure for RL-based query augmentation can be summarized as Algorithm˜1. This formulation allows training a rewriter even when the retrieval system is a black-box, such as an API-based service, emphasizing the connection between query generation and retrieval performance.